# Rectification Reimagined: A Unified Mamba Model for Image Correction and Rectangling with Prompts

### v1

@@ Abstract
Figure 1: (a) Pragmatic tasks for a smartphone. We consider four detailed tasks for image correction and rectangling, which are closely related to two types of common cameras on a mainstream mobile phone. Portrait correction ($\mathcal{T}_{1}$) and rectified wide-angle image rectangling ($\mathcal{T}_{2}$) often need to process pictures taken with a wide-angle lens. Stitched image rectangling ($\mathcal{T}_{3}$) and rotation correction ($\mathcal{T}_{4}$) are two practical tasks for daily life. (b) Camera distortion models and flow visualization for different tasks. On a normalized plane, the image of the point $P$ is $p_{d}$ whereas it would be $p$ without distortion following a pinhole camera model (kannala2006generic). The optical flows of backward distortions are generated through the RAFT (teed2020raft), which is the one of the most powerful tools for optical flow estimation. (c) Multi-task types. (1) four-to-four. Previous studies often use four disparate models to accomplish four tasks separately. (2) four-by-one. four tasks are achieved by one model structure but sharing different even counteractive network weights. (3) four-in-one. four tasks are employed in one model which can handle diverse tasks at the same time. four-by-one and four-in-one are all task-agnostic models. We obtain the four-by-one and four-in-one in this paper.
@@ Introduction
With the rapid development of photography system, there has been a surging interest in low-level tasks, which play pivotal roles in enhancing image quality. Image correction usually aims at estimating and correcting distortions (li2019blind; zhao2018distortion), and has a very wide range of applications (zhou2018gridface; zhuang2019learning; xue2019learning; zhan2019esir). Image rectangling is a special type of post-processing task (he2013rectangling) to acquire user-friendly rectangular images for photos with irregular boundaries while maintaining high content fidelity. Previous studies have addressed these problems by devising diverse specific architectures to learn different types of processing. For instance, tan2021practical proposed a cascaded network and progressively correct portrait distortion. nie2022deep optimized a mesh model for stitched image rectangling. Nearly all of these methods concentrate on a single task and depend on task-specific networks. Besides, we note that these practical tasks may be required within one smartphone device and associated with specific cameras or lenses, as depicted in Fig. 1a. The former indicates that identifying a unified task-agnostic model could be highly beneficial for edge devices with limited computation capabilities and memory resources. The latter suggests that potential distortion relationships may exist in other tasks beyond the portrait and rotation correction tasks when considering all tasks as one general rectification task.
@@
Image Correction. Portrait distortion is a common issue encountered in smartphone photography. shih2019distortion proposed an algorithm based on mesh optimization to restore the wrapped content. tan2021practical introduced the first end-to-end method based on deep learning and provided a supervised dataset and zhu2022semi proposed a semi-supervised method to further improve results. Rotation distortion is caused by arbitrary rotation angle. Direct rotation requires angle-prior and also destroys the rectangular boundaries. Some previous methods preserve image structure (edge or line) to ensure the naturalness of image content (he2013content; von2008lsd). nie2023deep proposed a mesh transformation network to correct content and boundaries. However, all these networks need task-specific designs.
@@
Image Rectangling. To obtain rectangular images after stitching, simple inner rectangle clipping leads to the loss of pixel semantics, while certain image inpainting or completion methods (suvorov2022resolution; teterwak2019boundless; liao2021towards) pose the risk of introducing uncertain information. he2013rectangling first introduced a content-aware algorithm that utilizes an energy function to optimize local mesh deformation. Recent DNN-based methods (nie2022deep; zhou2024recdiffusion; qiu2024remote) estimated the wrapping meshes or motion fields to handle non-rectangular borders. liao2023recrecnet constructed a win-win representation for rectified wide-angle image rectangling. In this work, we regard these tasks as a unified rectification task.
@@
Multi-Task Learning. Multi-task learning methods can be roughly divided into task balancing and multi-task architecture. Tasking balancing (kendall2018multi; guo2018dynamic) methods focus on weights on task and gradient conflicts by re-weighting the loss or manipulating the gradient. This type of method can successfully address multi-task learning for similar problems (liao2024mowa), such as simultaneous image denoising and deblurring (zamir2021multi; potlapalli2024promptir). However, it tends to underperform when there is a significant disparity between different tasks. Besides, some other approaches (bragman2019stochastic; ruder2019latent; gao2019nddr) propose the advanced architecture for parameter sharing. gao2019nddr employed different decoders for different tasks, while ruder2019latent proposed a gating mechanism as soft parameters sharing between different task networks. We try to mitigate task competition and achieve the four-in-one in this paper.
@@
Figure 2: (a) Framework of our Unified Rectification. It mainly consists of a (b) deformation module (DM) and a (d) restoration module (RM), which are trained simultaneously. An image $X^{i}_{0}$ from the image set of task $\mathcal{T}_{i}$ and a corresponding visual prompt $M^{i}_{0}$ indicating which task to perform are set as input for UniRect, which yields the final result $X^{i}_{R}$. In deformation module (ignoring our residual progressive setting for simplicity), the (c) control point predictor $\mathcal{C}$ can predict the locations of a set of control points, i.e. $c$ , with which the grid generator $\mathcal{G}$ produces a sampling grid $\mathcal{P}$ by Eq. 5. The sampler $\mathcal{S}$ then samples from $X^{i}_{0}$ and $M^{i}_{0}$ with the restriction of $\mathcal{P}$, resulting in the rectified image $X^{i}_{D}$ and its new prompt $M^{i}_{D}$. For some tasks without boundary changes like $\mathcal{T}_{1}$ and $\mathcal{T}_{4}$, $M^{i}_{D}$ will be changed into the all-one matrix. (e) Mamba block is applied in the control point predictor to obtain geometric information like borders by its scan characteristic (gu2021combining; gu2023mamba). RM is composed of residual mamba blocks (RMBs), which effectively captures global connections for restoration. (f) Rectification for all tasks. Our model treats all tasks as rectification tasks and subsequently incorporates them into a unified framework.
@@
1) For portrait distortion in $\mathcal{T}_{1}^{-1}$, it has been fully studied by previous researchers and the Kanala-Brandt model (kannala2006generic) is widely used to approximate the distortion of a wide-angle lens: $r_{d}=\sum_{j=1}^{N}{k_{j}{\theta^{2j-1}}},N=1,2,3,\cdots$, where $\theta$ denotes the angle between the incoming ray and the optical axis and $k_{j}$ is the distortion parameter. 2) For wide-angle image rectified distortion in $\mathcal{T}_{2}^{-1}$, we observe that its flow appears a radial distribution as indicated by the red arrow in Fig. 1b.2 and a traditional Brown-Conrady model (weng1992camera) in introduced here: $r_{d}=\sum_{j=1}^{N}{k_{j}{r^{2j-1}}},N=1,2,3,\cdots$. 3) The stitched distortion in $\mathcal{T}^{-1}_{3}$ is formed by supposing that the rectangular image is transmuted into a image with irregular boundaries through an amorphous and vibrant lens. We discover that the flow in Fig. 1b.3 presents an approximate radial distribution around a particular point $O_{2}^{\prime}$, which is affected by irregular borders. Thus, we biuld a variant of Brown-Conrady model to describe this phenomenon:
@@
where $\Phi$ denotes the transformation and $\lambda$ is a weight to balance two terms. The only closed-form solution (kent1994link) of Eq. 4 can be derived as
@@
Network Structures. The structure of Control Point Predictor $\mathcal{C}_{0,1}$ is shown in Fig. 2c. The image $X^{i}_{0,1}$ and its prompt $M^{i}_{0,1}$ are concatenated and fed into $\mathcal{C}_{0,1}$ to generate the deviation locations $\Delta c_{0,1}\in\mathbb{R}^{B\times N_{c}\times 2}$ of control points $P^{\prime}$ in $X^{i}_{0,1}$. Inspired by recent burgeoning development of Mamba (gu2021combining; gu2023mamba), a mamba block is introduced in the latent space to scan the geometric information and handle insidious long-range dependencies of distortions. The structure of Restoration Module is shown in Fig. 2d, which is mainly composed of four Residual Mamba Blocks (RMBs). In practice, each RMB has $32$ channels in convolution layers and a same mamba block is in the core of RMB to capture non-local connections effectively like many vision tasks (liu2024vmamba; zhu2024vision; guo2025mambair). $X^{i}_{D}$ and $M^{i}_{D}$ first pass through a partial convolution layer (Liu2018ImageIF) as there may be some irregular borders in $X^{i}_{D}$ for boundary-changing task like $\mathcal{T}_{2},\mathcal{T}_{3},\mathcal{T}_{3}$ due to the inaccurate estimation of the positions of the control points in the previous module. For $\mathcal{T}_{1},\mathcal{T}_{4}$, the partial convolution is equivalent to a standard convolution layer since the $M^{i}_{D}$ is set to all-one matrix.
@@
where $\partial M$ is the zero level set. $M_{in}$ and $M_{out}$ represent the inside area and outside area for $M$. We also use the appearance loss $\mathcal{L}_{a}$ (i.e. L1 loss), line and shape penalty $\mathcal{L}_{p}$ (nie2022deep; liao2023recrecnet) on the mesh formed by $P^{\prime}$ to prevent excessive rectification, and the gradient loss $\mathcal{L}_{g}$ to align local textures. Finally, the all loss $\mathcal{L}_{DM}$ can be expressed by
@@
where $j=0$ for $X_{1}^{i}$ and $j=1$ for $X_{D}^{i}$. $\gamma,\alpha_{1},\alpha_{2},\alpha_{3}$ are loss weights. Note that we only apply $\mathcal{L}_{a}$ for $\mathcal{T}_{1},\mathcal{T}_{4}$ since these two tasks have no border changes during rectification. For restoration module, we only use appearance loss and perceptual loss like some super-resolution works (Wang2018ESRGANES; johnson2016perceptual).
@@
where $G(\cdot)$ is a gating network and $E_{j}$ is the $j$-th expert network (i.e. our UniRect). This gating network can apportion the specific input task among $E_{j}$ with a $5$-dimension vector weights. Specifically, to reduce computation, $G$ with network weights $W_{G}$ has a additional top-$k$ operator, enforcing the invalid value to be zero (shazeer2017outrageously), i.e.
@@
Portrait correction $\mathcal{T}_{1}$ shih2019distortion 66.143 97.253 - -
tan2021practical 66.784 97.490 - -
zhu2022semi 66.825 97.491 7.082 8.082
@@
Rectified wide- angle image rectangling $\mathcal{T}_{2}$ ROP (liao2021towards) 13.90 0.3516 - -
RecRecNet (liao2023recrecnet) 18.68 0.5450 19.01 0.1136
@@
Stitched image rectangling $\mathcal{T}_{3}$ he2013rectangling 14.70 0.3775 38.19 0.2846
nie2022deep 21.28 0.7141 21.77 0.1557
MOWA (liao2024mowa) 20.42 0.6307 - -
@@
Rotation correction $\mathcal{T}_{4}$ he2013content 21.69 0.6460 8.51 0.2120
DRC (nie2023deep) 21.02 0.6280 7.12 0.2050
CoupledTPS (nie2024semisupervised) 22.29 0.6790 7.90 0.1970
@@
We conduct experiments on public representative benchmarks including portrait correction dataset (tan2021practical), rectified wide-angle image rectangling dataset (liao2023recrecnet), stitched image rectangling dataset (nie2022deep), and rotation correction dataset (nie2023deep).
@@
Our networks are based on the PyTorch framework with four NVIDIA Tesla V100 GPUs. For single-task learning, it takes one to three days to train our model on each dataset with a batch size of $4$ and 200 epochs. For multi-task learning, it takes about $7$ days to train on four datasets jointly with a bacth size of $10$ and 200 epochs. We adopt the poly policy of the learning rate, the factor of which is $0.96$, and the initial learning rate is 1e51e-5 for rotation correction and 1e41e-4 for the rest tasks. Adam (kingma2014adam) optimizer is selected and the weight decay is 1e51e-5. For RP-TPS, the number of control points is set to $12\times 10$. For Eq. 9, we set $\gamma,\alpha_{1},\alpha_{2},\alpha_{3}$ as 0.9,1e2,1.0,1e20.9,1e-2,1.0,1e-2 respectively. For SMoEs, we utilize a simple ResNet18 (he2016deep) as the backbone of $G$. For computation complexity, the amounts of parameters (Params), floating point operations (FLOPs), and frames per second (FPS) are 357.9M, 62.98G, and 35.8 respectively. More analysis for computations can be seen in the supplementary files.
@@
We adopt the evaluation setting from the previous studies, utilizing the PSNR, SSIM (Wang2004ImageQA), FID (heusel2017gans) and LPIPS (Zhang2018TheUE) to assess these methods for $\mathcal{T}_{2},\mathcal{T}_{3},\mathcal{T}_{4}$. The first two are objective metrics while the following are image perceptual metrics. For portrait correction, Line Straightness Metric (LineACC) and Shape Congruence Metric (ShapeACC) are suggested in (tan2021practical) to measure the correction results of salient lines and faces. Besides, many studies (zhang2023gpt; wu2024comprehensive; xie2024modification) have proposed exploiting MLLMs to assess the quality of unlabeled images, which has led to remarkable progress in the field. Owing to the limitations  [^1] In the test set of this dataset, only the coordinates of the corresponding points in the reference and input distorted images are provided for metric calculation. Nevertheless, in our approach, because of the characteristics of the TPS model (irreversibility), it is arduous to acquire the precise coordinates of the key points on the distorted image after rectification, as depicted in Fig. 2f. Hence, we resort to an estimation method, resulting in a substantial reduction in the metrics of us. of these two evaluation metrics, we further utilize Multimodal Large Language Models (MLLMs)-based metrics, namely LineACC-LLM and ShapeACC-LLM. By capitalizing on the potent capabilities of MLLMs, our objective is to carry out an assessment of image quality in terms of line straightness and shape congruence. Further details about these metrics can be found in the supplementary files.

### v2

@@ Abstract
Figure 1: (a) Pragmatic tasks for a smartphone. We consider four detailed tasks for image correction and rectangling, which are closely related to two types of common cameras on a mainstream mobile phone. Portrait correction ($\mathcal{T}_{1}$) and rectified wide-angle image rectangling ($\mathcal{T}_{2}$) often need to process pictures taken with a wide-angle lens. Stitched image rectangling ($\mathcal{T}_{3}$) and rotation correction ($\mathcal{T}_{4}$) are two practical tasks for daily life. (b) Camera distortion models and flow visualization for different tasks. On a normalized plane, the image of the point $P$ is $p_{d}$ whereas it would be $p$ without distortion following a pinhole camera model (Kannala and Brandt 2006). The optical flows of backward distortions are generated through the RAFT (Teed and Deng 2020), which is the one of the most powerful tools for optical flow estimation. (c) Multi-task types. (1) four-to-four. Previous studies often use four disparate models to accomplish four tasks separately. (2) four-by-one. four tasks are achieved by one model structure but sharing different even counteractive network weights. (3) four-in-one. four tasks are employed in one model which can handle diverse tasks at the same time. four-by-one and four-in-one are all task-agnostic models. We obtain the four-by-one and four-in-one in this paper.
@@ Introduction
With the rapid development of photography system, there has been a surging interest in low-level tasks, which play pivotal roles in enhancing image quality. Image correction usually aims at estimating and correcting distortions (Li et al. 2019; Zhao et al. 2018), and has a very wide range of applications (Zhou et al. 2018; Zhuang et al. 2019; Xue et al. 2019; Zhan and Lu 2019). Image rectangling is a special type of post-processing task (He et al. 2013b) to acquire user-friendly rectangular images for photos with irregular boundaries while maintaining high content fidelity. Previous studies have addressed these problems by devising diverse specific architectures to learn different types of processing. For instance, Tan et al. (2021) proposed a cascaded network and progressively correct portrait distortion. Nie et al. (2022) optimized a mesh model for stitched image rectangling. Nearly all of these methods concentrate on a single task and depend on task-specific networks. Besides, we note that these practical tasks may be required within one smartphone device and associated with specific cameras or lenses, as depicted in Fig. 1a. The former indicates that identifying a unified task-agnostic model could be highly beneficial for edge devices with limited computation capabilities and memory resources. The latter suggests that potential distortion relationships may exist in other tasks beyond the portrait and rotation correction tasks when considering all tasks as one general rectification task.
@@
Image Correction. Portrait distortion is a common issue encountered in smartphone photography. Shih et al. (2019) proposed an algorithm based on mesh optimization to restore the wrapped content. Tan et al. (2021) introduced the first end-to-end method based on deep learning and provided a supervised dataset and Zhu et al. (2022) proposed a semi-supervised method to further improve results. Rotation distortion is caused by arbitrary rotation angle. Direct rotation requires angle-prior and also destroys the rectangular boundaries. Some previous methods preserve image structure (edge or line) to ensure the naturalness of image content (He et al. 2013a; Von Gioi et al. 2008). Nie et al. (2023) proposed a mesh transformation network to correct content and boundaries. However, all these networks need task-specific designs.
@@
Image Rectangling. To obtain rectangular images after stitching, simple inner rectangle clipping leads to the loss of pixel semantics, while certain image inpainting or completion methods (Suvorov et al. 2022; Teterwak et al. 2019; Liao et al. 2021a) pose the risk of introducing uncertain information. He et al. (2013b) first introduced a content-aware algorithm that utilizes an energy function to optimize local mesh deformation. Recent DNN-based methods (Nie et al. 2022; Zhou et al. 2024; Qiu et al. 2024) estimated the wrapping meshes or motion fields to handle non-rectangular borders. Liao et al. (2023) constructed a win-win representation for rectified wide-angle image rectangling. In this work, we regard these tasks as a unified rectification task.
@@
Multi-Task Learning. Multi-task learning methods can be roughly divided into task balancing and multi-task architecture. Tasking balancing (Kendall et al. 2018; Guo et al. 2018) methods focus on weights on task and gradient conflicts by re-weighting the loss or manipulating the gradient. This type of method can successfully address multi-task learning for similar problems (Liao et al. 2025), such as simultaneous image denoising and deblurring (Zamir et al. 2021; Potlapalli et al. 2024). However, it tends to underperform when there is a significant disparity between different tasks. Besides, some other approaches (Bragman et al. 2019; Ruder et al. 2019; Gao et al. 2019) propose the advanced architecture for parameter sharing. Gao et al. (2019) employed different decoders for different tasks, while Ruder et al. (2019) proposed a gating mechanism as soft parameters sharing between different task networks. We try to mitigate task competition and achieve the four-in-one in this paper.
@@
Figure 2: (a) Framework of our Unified Rectification. It mainly consists of a (b) deformation module (DM) and a (d) restoration module (RM), which are trained simultaneously. An image $X^{i}_{0}$ from the image set of task $\mathcal{T}_{i}$ and a corresponding visual prompt $M^{i}_{0}$ indicating which task to perform are set as input for UniRect, which yields the final result $X^{i}_{R}$. In deformation module (ignoring our residual progressive setting for simplicity), the (c) control point predictor $\mathcal{C}$ can predict the locations of a set of control points, i.e. $c$ , with which the grid generator $\mathcal{G}$ produces a sampling grid $\mathcal{P}$ by Eq. 5. The sampler $\mathcal{S}$ then samples from $X^{i}_{0}$ and $M^{i}_{0}$ with the restriction of $\mathcal{P}$, resulting in the rectified image $X^{i}_{D}$ and its new prompt $M^{i}_{D}$. For some tasks without boundary changes like $\mathcal{T}_{1}$ and $\mathcal{T}_{4}$, $M^{i}_{D}$ will be changed into the all-one matrix. (e) Mamba block is applied in the control point predictor to obtain geometric information like borders by its scan characteristic (Gu et al. 2021; Gu and Dao 2023). RM is composed of residual mamba blocks (RMBs), which effectively captures global connections for restoration. (f) Rectification for all tasks. Our model treats all tasks as rectification tasks and subsequently incorporates them into a unified framework.
@@
1) For portrait distortion in $\mathcal{T}_{1}^{-1}$, it has been fully studied by previous researchers and the Kanala-Brandt model (Kannala and Brandt 2006) is widely used to approximate the distortion of a wide-angle lens: $r_{d}=\sum_{j=1}^{N}{k_{j}{\theta^{2j-1}}},N=1,2,3,\cdots$, where $\theta$ denotes the angle between the incoming ray and the optical axis and $k_{j}$ is the distortion parameter. 2) For wide-angle image rectified distortion in $\mathcal{T}_{2}^{-1}$, we observe that its flow appears a radial distribution as indicated by the red arrow in Fig. 1b.2 and a traditional Brown-Conrady model (Weng et al. 1992) in introduced here: $r_{d}=\sum_{j=1}^{N}{k_{j}{r^{2j-1}}},N=1,2,3,\cdots$. 3) The stitched distortion in $\mathcal{T}^{-1}_{3}$ is formed by supposing that the rectangular image is transmuted into a image with irregular boundaries through an amorphous and vibrant lens. We discover that the flow in Fig. 1b.3 presents an approximate radial distribution around a particular point $O_{2}^{\prime}$, which is affected by irregular borders. Thus, we biuld a variant of Brown-Conrady model to describe this phenomenon:
@@
where $\Phi$ denotes the transformation and $\lambda$ is a weight to balance two terms. The only closed-form solution (Kent and Mardia 1994) of Eq. 4 can be derived as
@@
Network Structures. The structure of Control Point Predictor $\mathcal{C}_{0,1}$ is shown in Fig. 2c. The image $X^{i}_{0,1}$ and its prompt $M^{i}_{0,1}$ are concatenated and fed into $\mathcal{C}_{0,1}$ to generate the deviation locations $\Delta c_{0,1}\in\mathbb{R}^{B\times N_{c}\times 2}$ of control points $P^{\prime}$ in $X^{i}_{0,1}$. Inspired by recent burgeoning development of Mamba (Gu et al. 2021; Gu and Dao 2023), a mamba block is introduced in the latent space to scan the geometric information and handle insidious long-range dependencies of distortions. The structure of Restoration Module is shown in Fig. 2d, which is mainly composed of four Residual Mamba Blocks (RMBs). In practice, each RMB has $32$ channels in convolution layers and a same mamba block is in the core of RMB to capture non-local connections effectively like many vision tasks (Liu et al. 2024; Zhu et al. 2024; Guo et al. 2025). $X^{i}_{D}$ and $M^{i}_{D}$ first pass through a partial convolution layer (Liu et al. 2018) as there may be some irregular borders in $X^{i}_{D}$ for boundary-changing task like $\mathcal{T}_{2},\mathcal{T}_{3},\mathcal{T}_{3}$ due to the inaccurate estimation of the positions of the control points in the previous module. For $\mathcal{T}_{1},\mathcal{T}_{4}$, the partial convolution is equivalent to a standard convolution layer since the $M^{i}_{D}$ is set to all-one matrix.
@@
where $\partial M$ is the zero level set. $M_{in}$ and $M_{out}$ represent the inside area and outside area for $M$. We also use the appearance loss $\mathcal{L}_{a}$ (i.e. L1 loss), line and shape penalty $\mathcal{L}_{p}$ (Nie et al. 2022; Liao et al. 2023) on the mesh formed by $P^{\prime}$ to prevent excessive rectification, and the gradient loss $\mathcal{L}_{g}$ to align local textures. Finally, the all loss $\mathcal{L}_{DM}$ can be expressed by
@@
where $j=0$ for $X_{1}^{i}$ and $j=1$ for $X_{D}^{i}$. $\gamma,\alpha_{1},\alpha_{2},\alpha_{3}$ are loss weights. Note that we only apply $\mathcal{L}_{a}$ for $\mathcal{T}_{1},\mathcal{T}_{4}$ since these two tasks have no border changes during rectification. For restoration module, we only use appearance loss and perceptual loss like some super-resolution works (Wang et al. 2018; Johnson et al. 2016).
@@
where $G(\cdot)$ is a gating network and $E_{j}$ is the $j$-th expert network (i.e. our UniRect). This gating network can apportion the specific input task among $E_{j}$ with a $5$-dimension vector weights. Specifically, to reduce computation, $G$ with network weights $W_{G}$ has a additional top-$k$ operator, enforcing the invalid value to be zero (Shazeer et al. 2017), i.e.
@@
Portrait correction $\mathcal{T}_{1}$ Shih et al. (2019) 66.143 97.253 - -
Tan et al. (2021) 66.784 97.490 - -
Zhu et al. (2022) 66.825 97.491 7.082 8.082
@@
Rectified wide- angle image rectangling $\mathcal{T}_{2}$ ROP (Liao et al. 2021a) 13.90 0.3516 - -
RecRecNet (Liao et al. 2023) 18.68 0.5450 19.01 0.1136
@@
Stitched image rectangling $\mathcal{T}_{3}$ He et al. (2013b) 14.70 0.3775 38.19 0.2846
Nie et al. (2022) 21.28 0.7141 21.77 0.1557
MOWA (Liao et al. 2025) 20.42 0.6307 - -
@@
Rotation correction $\mathcal{T}_{4}$ He et al. (2013a) 21.69 0.6460 8.51 0.2120
DRC (Nie et al. 2023) 21.02 0.6280 7.12 0.2050
CoupledTPS (Nie et al. 2024) 22.29 0.6790 7.90 0.1970
@@
We conduct experiments on public representative benchmarks including portrait correction dataset (Tan et al. 2021), rectified wide-angle image rectangling dataset (Liao et al. 2023), stitched image rectangling dataset (Nie et al. 2022), and rotation correction dataset (Nie et al. 2023).
@@
Our networks are based on the PyTorch framework with four NVIDIA Tesla V100 GPUs. For single-task learning, it takes one to three days to train our model on each dataset with a batch size of $4$ and 200 epochs. For multi-task learning, it takes about $7$ days to train on four datasets jointly with a bacth size of $10$ and 200 epochs. We adopt the poly policy of the learning rate, the factor of which is $0.96$, and the initial learning rate is 1e51e-5 for rotation correction and 1e41e-4 for the rest tasks. Adam (Kingma and Ba 2014) optimizer is selected and the weight decay is 1e51e-5. For RP-TPS, the number of control points is set to $12\times 10$. For Eq. 9, we set $\gamma,\alpha_{1},\alpha_{2},\alpha_{3}$ as 0.9,1e2,1.0,1e20.9,1e-2,1.0,1e-2 respectively. For SMoEs, we utilize a simple ResNet18 (He et al. 2016) as the backbone of $G$. For computation complexity, the amounts of parameters (Params), floating point operations (FLOPs), and frames per second (FPS) are 357.9M, 62.98G, and 35.8 respectively. More analysis for computations can be seen in the supplementary files.
@@
We adopt the evaluation setting from the previous studies, utilizing the PSNR, SSIM (Wang et al. 2004), FID (Heusel et al. 2017) and LPIPS (Zhang et al. 2018) to assess these methods for $\mathcal{T}_{2},\mathcal{T}_{3},\mathcal{T}_{4}$. The first two are objective metrics while the following are image perceptual metrics. For portrait correction, Line Straightness Metric (LineACC) and Shape Congruence Metric (ShapeACC) are suggested in (Tan et al. 2021) to measure the correction results of salient lines and faces. Besides, many studies (Zhang et al. 2023; Wu et al. 2024; Xie et al. 2024) have proposed exploiting MLLMs to assess the quality of unlabeled images, which has led to remarkable progress in the field. Owing to the limitations  [^1] In the test set of this dataset, only the coordinates of the corresponding points in the reference and input distorted images are provided for metric calculation. Nevertheless, in our approach, because of the characteristics of the TPS model (irreversibility), it is arduous to acquire the precise coordinates of the key points on the distorted image after rectification, as depicted in Fig. 2f. Hence, we resort to an estimation method, resulting in a substantial reduction in the metrics of us. of these two evaluation metrics, we further utilize Multimodal Large Language Models (MLLMs)-based metrics, namely LineACC-LLM and ShapeACC-LLM. By capitalizing on the potent capabilities of MLLMs, our objective is to carry out an assessment of image quality in terms of line straightness and shape congruence. Further details about these metrics can be found in the supplementary files.